OpenStudy (raveninneverland):

Give the range of x - 2y = 0 for the domain {-1, 0, 4} How would I find the range??

OpenStudy (anonymous):

Domain is basically what you plug in the x, and range is the y So, let's go ahead and first get the equation ready so we can plug it in quickly. x-2y=0 x=2y x/2 = y Now let's put in the three domains given into the x -1/2 for -1 0/2 = 0 for 0 4/2 = 2 for 4

OpenStudy (raveninneverland):

Alright, I think I've caught on now! So, just divide 2 in the domain, and I'll get the range?

OpenStudy (anonymous):

Yes, for this particular equation. I just moved the equation around to make it easier to plug in, so you would just have to divide by 2. However, if there's a DIFFERENT equation in a future problem, you can just plug it into the x directly and solve for y, which would be the range.

OpenStudy (raveninneverland):

Okay, this has definitely helped! Thank you. :)